# AraFinNews: Arabic Financial Summarisation with Domain-Adapted LLMs

### v2

@@ Abstract
This paper investigates the impact of domain specificity on abstractive summarisation of Arabic financial texts using large language models (LLMs). We introduce AraFinNews, the largest publicly available Arabic financial news dataset to date, comprising 212,500 article–headline pairs spanning nearly a decade of reporting from October 2015 to July 2025. Designed as the Arabic equivalent of major English summarisation corpora such as CNN/DailyMail, AraFinNews provides a robust benchmark for evaluating domain-specific language understanding and generation in financial contexts. Using this resource, we evaluate transformer-based models—including mT5, AraT5, and the domain-adapted FinAraT5—to examine how financial-domain pretraining influences accuracy, numerical reliability, and stylistic alignment with professional reporting. Experimental results show that domain-adapted models generate more coherent summaries, particularly in handling quantitative and entity-centric information. The findings highlight the importance of domain-specific adaptation for improving narrative fluency in Arabic financial summarisation. The dataset is freely available for non-commercial research athttps://github.com/ArabicNLP-uk/AraFinNews
@@ I Introduction
Arabic, spoken by over 300 million people across more than twenty countries [statista2024], holds central importance in religious, political, and cultural communication throughout the Arab world. Despite its global prominence, Arabic remains underrepresented in Natural Language Processing (NLP). While initiatives such as ACL SIGARAB [^1] https://www.sigarab.org/ have improved visibility, progress in building robust, domain-specific Arabic language models and datasets continues to lag behind high-resource languages such as English and Chinese [el2010using]. This gap stems not only from data scarcity but also from the language’s structural complexity: Arabic is diglossic, encompassing Classical Arabic, Modern Standard Arabic (MSA), and diverse dialects that differ lexically, morphologically, and syntactically. Moreover, its right-to-left script, absence of capitalisation, and omission of diacritics—short vowel markers that disambiguate homographs—pose challenges for tokenisation, parsing, and semantic interpretation [el-haj-rayson-2016-osman].
Recent advances in Arabic NLP through models such as AraBERT [antoun_arabert_2021], ARBERT, MARBERT [abdul-mageed_arbert_2021], and generative models like AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022] have substantially improved general text understanding. Yet, these models are trained on broad web or news data, lacking exposure to specialised financial or technical language. In contrast, English NLP has benefited from domain-adapted architectures such as FinBERT [yang2020finbert] and BioBERT [lee2020biobert], which deliver notable performance gains in financial and biomedical applications. The absence of comparable domain adaptation for Arabic highlights the need for targeted pretraining strategies to capture the linguistic and conceptual nuances of financial discourse.
The Arab financial sector has rapidly expanded over the past two decades, with markets such as Tadawul [^2] https://www.saudiexchange.sa, the Dubai Financial Market (DFM) [^3] https://www.dfm.ae, and the Abu Dhabi Securities Exchange (ADX) [^4] https://www.adx.ae now ranking among the world’s most active [zmandar2021joint]. This growth has generated a surge of Arabic financial narratives—ranging from corporate disclosures and regulatory filings to market analyses and news reports—typically long, information-dense, and stylistically formal, making them ideal for automatic summarisation [zmandar2021multilingual]. Yet most Arabic summarisation systems focus on general news, leaving financial text underexplored despite its demand for accuracy, numerical precision, and stylistic clarity [el2020financial, zmandar2021financial, zavitsanos2023financial].
@@ II Related Work
Text summarisation has long been a key task in Natural Language Processing (NLP), particularly in journalism and information retrieval. Early systems were largely extractive, relying on sentence ranking and statistical heuristics such as term frequency, cue words, and sentence position [filippova2009company, berger2000ocelot]. While these methods captured surface salience, they lacked true abstraction and often produced redundant or disjointed summaries.
@@
The emergence of neural encoder–decoder architectures marked a significant shift. Models based on recurrent neural networks (RNNs) with attention mechanisms introduced the ability to generate novel, coherent text rather than copy input segments verbatim. One of the first neural systems capable of producing fluent news headlines using LSTM-based architectures was demonstrated in [lopyrev2015generating]. Later developments integrated coverage mechanisms and reinforcement learning to address factual consistency and redundancy [see-etal-2017-get, cheng-etal-2016-neural].
@@
Transformer-based models, notably BART and T5, further revolutionised abstractive summarisation. By treating all NLP tasks as text-to-text transformations, these architectures achieved strong generalisation across domains and languages [raffel_exploring_2020]. Their multilingual variants, such as mT5 [xue-etal-2021-mt5], extended these capabilities to low-resource languages through cross-lingual transfer. While effective in general domains, these models often struggle in specialised contexts—such as finance—where domain-specific terminology, numerical reasoning, and formal discourse require more focused adaptation.
@@
In recent years, Arabic NLP has seen rapid progress due to the introduction of large-scale pretrained language models. Encoder-based architectures like AraBERT [antoun_arabert_2021], ARBERT and MARBERT[abdul-mageed_arbert_2021], have become benchmarks for classification and sentiment analysis tasks, addressing key challenges such as morphology and dialectal variation. Generative architectures including AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022], and more recent efforts such as Jasmine [abdul2023jasmine], have extended these capabilities to summarisation and text generation.
@@
A growing body of work now focuses on Arabic abstractive summarisation. [kahla2023fine] demonstrated the effectiveness of fine-tuning multilingual transformers for Arabic news summarisation, while [morsi2025arabic] proposed transformer-based architectures that improved both coherence and fluency over traditional models. Other work has explored integrating linguistic knowledge or auxiliary tasks: for instance, [essa2025enhanced] combined named entity recognition with summarisation to better preserve factual content, and [elsaid2025hybrid] developed a hybrid RNN–Transformer model to improve grammatical consistency. These advances collectively demonstrate the viability of generative models for Arabic, though the majority of research remains limited to general news rather than specialised professional domains.
Considerable advances have been made in Arabic language modelling with the introduction of pretrained systems such as AraBERT [antoun_arabert_2021], ARBERT [abdul-mageed_arbert_2021], MARBERT, and generative variants like AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022]. However, these remain largely general-purpose, drawing on broad web and news corpora rather than domain-specific text. In English, models like FinBERT [yang2020finbert] and BioBERT [lee2020biobert] have established the effectiveness of domain-oriented pretraining, yet equivalent efforts for Arabic are still limited. Existing Arabic models show strong performance in generic understanding tasks but struggle with the precision and structure required for financial narrative summarisation.
@@ II-A Datasets for Arabic Summarisation and Financial Corpora
Research in Arabic summarisation has been supported by a range of corpora covering both extractive and abstractive paradigms. Early efforts such as Kalimat [ElHaj2013KALIMATAM], a 20,000-article Omani newspaper corpus, and the Essex Arabic Summaries Corpus (EASC) [el2010using], which includes 153 news articles and 765 human-written summaries, provided the foundation for extractive summarisation. More recent multilingual datasets have enabled progress in neural and cross-lingual approaches. The Arabic Gigaword corpus [Parker2019gigawords], although large and widely used, lacks aligned document–summary pairs and remains commercially restricted, limiting its direct use for supervised learning. By contrast, XLSum [hasan-etal-2021-xl] and WikiLingua [ladhak-etal-2020-wikilingua] introduced openly available abstractive benchmarks with approximately 40,000 and 21,000 Arabic examples respectively—though their content and style differ from journalistic and financial text. Additional corpora such as SumArabic, which offers over 84,000 abstractive summaries, and HASD, containing 43,000 hybrid samples, have expanded the scale of available data but still show inconsistencies in length and domain coverage.
Despite these advances, none of the existing datasets focus specifically on financial news, leaving a gap for domain-specific summarisation resources. The AraFinNews dataset introduced in this paper directly addresses this gap, offering over 200,000 high-quality Arabic financial article–headline pairs for research in summarisation and broader financial NLP.
@@
Early Arabic summarisation research applied extractive techniques including TF–IDF ranking, lexical scoring, and graph-based centrality measures [El-Haj-2011, Douzidia2004LakhasAA, el2013using, koulali2013arabic]. More recently, multilingual models such as mBART and mT5 have been fine-tuned for Arabic summarisation [kahla-etal-2021-cross], achieving improvements in lexical diversity and fluency.
@@
Unlike English, Arabic still lacks large, well-curated corpora for specialised domains. Existing resources such as BORSAH [ALSHAHRANI18] and the Arabic Business and Management Corpus (ABMC) [^5] https://elhaj.uk/corpora.html provide valuable text collections in finance and economics, but they were primarily designed for sentiment analysis or classification rather than summarisation. As a result, progress in Arabic abstractive summarisation—particularly within financial contexts—has been constrained by the absence of domain-focused, large-scale, high-quality datasets.
@@
The limited intersection of Arabic summarisation and financial text processing highlights an important research opportunity: developing domain-adapted datasets and models capable of handling numerically dense, terminologically specific, and stylistically formal text. Such work is crucial for advancing the broader goal of creating practical, context-aware summarisation systems for professional Arabic applications.
@@ III AraFinNews: The Arabic Financial News Dataset
To address the lack of large, domain-specific resources for Arabic summarisation, we introduce AraFinNews, the largest publicly available Arabic financial news dataset to date. The corpus comprises 212,500 article–headline pairs sourced through Argaam.com, a major financial media platform in the Gulf region that aggregates material from several well-established outlets, including the Financial Times, Forbes, the World Gold Council and Bloomberg. Each headline serves as an abstractive summary of its corresponding article, reflecting the concise, information-rich style characteristic of professional Arabic financial journalism. The dataset spans corporate disclosures, market analyses and macroeconomic reports, offering a solid benchmark for abstractive summarisation as well as related NLP tasks such as named entity recognition, event extraction and sentiment analysis.
AraFinNews covers a continuous period of nearly ten years, from 8 October 2015 to 29 July 2025, capturing key economic cycles, market events, and policy developments across the Arab world. This long-term temporal span ensures the dataset reflects evolving financial language, trends, and reporting styles over time—making it a valuable resource for both synchronic and diachronic analyses of Arabic financial discourse.
@@ IV Financial Large Language Models for Arabic Summarisation
To address this gap, Arabic-adapted transformer models such as AraBERT [antoun2020arabert] and AraT5 [nagoudi_arat5_2022] have extended large-scale pretraining to Modern Standard Arabic. More recently, domain-specialised variants such as FinAraT5 [zmandar2023finarat5] have been developed to capture the unique stylistic and numerical features of financial discourse.
In this study, we employ FinAraT5777https://huggingface.co/drelhaj/FinAraT5 alongside multilingual and general-purpose Arabic transformers to perform abstractive summarisation on AraFinNews. These models are evaluated under a unified experimental setup designed to assess their ability to generate coherent financial headlines. The following subsections outline the model selection, fine-tuning configuration, and evaluation framework adopted in this work.
@@ IV-A Model Selection and Comparative Setup
All models adopt a text-to-text paradigm supporting multiple generation tasks through consistent input–output formatting. Each was fine-tuned on the AraFinNews corpus using identical hyperparameters to ensure comparability. Fine-tuning employed the Hugging Face Transformers library with the AdamW optimiser, a learning rate of $5\times 10^{-5}$ (with linear warm-up), an effective batch size of 32 achieved via gradient accumulation of four steps, and a maximum sequence length of 512 tokens. Evaluation assesses how domain adaptation influences summarisation performance.
@@ IV-B Arabic Financial Summarisation Pipeline
This study employs FinAraT5 [zmandar2023finarat5], a domain-adapted generative language model fine-tuned primarily for Arabic financial headline generation and abstractive summarisation. Built upon the AraT5 architecture [nagoudi_arat5_2022] and following the text-to-text paradigm established by T5 [raffel_exploring_2020], FinAraT5 benefits from continued pretraining on large-scale Arabic financial corpora, including the AraFinNews dataset introduced in Section III. This adaptation enables the model to internalise the specialised terminology, discourse structures, and stylistic conventions characteristic of financial journalism.
Within our summarisation pipeline, FinAraT5 acts as the core component for abstractive headline generation, transforming full-length news articles into information-rich summaries aligned with the linguistic and stylistic expectations of Modern Standard Arabic financial communication.
While FinAraT5 follows the same sequence-to-sequence framework as T5, its architecture and training are optimised for summarisation rather than broad multitask generalisation (Figure 3).
@@
FinAraT5 was pretrained using the span-corruption denoising objective introduced in the original T5 framework [raffel_exploring_2020], where contiguous spans of text are randomly masked and the model learns to reconstruct the missing content. This self-supervised approach enables large-scale learning without labelled data and effectively captures long-range contextual dependencies within text.
Following the methodology of AraT5 [nagoudi_arat5_2022], FinAraT5 extends this objective through continued pretraining on raw Arabic financial documents, including large-scale news sources such as AraFinNews. This additional phase allows the model to internalise financial terminology, numerical expressions, and the discourse structures typical of professional economic reporting. As illustrated in Figure 4, the span-corruption objective provides an efficient mechanism for leveraging vast amounts of unannotated financial text while maintaining contextual richness and linguistic fidelity.
@@
FinAraT5 was trained using the original TensorFlow implementation of T5888https://github.com/google-research/text-to-text-transfer-transformer and managed via the t5 library [^9] https://pypi.org/project/t5/. Although subsequent frameworks such as T5X improved TPU scalability and efficiency, FinAraT5 predates their release and employed Mesh TensorFlow for distributed multi-core training.
This pretraining strategy establishes a strong foundation for downstream generative tasks, particularly abstractive summarisation. The model’s continued exposure to Arabic financial text enhances its treatment of specialised vocabulary and high-frequency expressions such as “stocks, profits, indices, returns”, giving it a marked advantage over general-purpose Arabic models like AraT5 and mT5. The fine-tuning configuration used in this study is summarised in Table V.

### v3

@@ Abstract
We introduce AraFinNews, the largest publicly available Arabic financial news dataset to date, comprising 212,500 article–headline pairs spanning a decade of reporting from 2015 to 2025. Designed as an Arabic counterpart to major English summarisation corpora such as CNN/DailyMail, AraFinNews provides a realistic benchmark for evaluating domain-specific language understanding and generation in financial contexts. Using this resource, we investigate the impact of domain specificity on abstractive summarisation of Arabic financial texts with large language models (LLMs). In particular, we evaluate transformer-based models—mT5, AraT5, and the domain-adapted FinAraT5—to examine how financial-domain pretraining influences accuracy, numerical reliability, and stylistic alignment with professional reporting. Experimental results show that domain-adapted models generate more coherent summaries, especially in their handling of quantitative and entity-centric information. These findings highlight the importance of domain-specific adaptation for improving narrative fluency in Arabic financial summarisation. The dataset is freely available for non-commercial research at https://github.com/ArabicNLP-uk/AraFinNews.
@@ I Introduction
Arabic, spoken by more than 300 million people across over twenty countries [statista2024], plays a central role in religious, political, and cultural communication throughout the Arab world. Yet, despite its global significance, the language remains underrepresented in Natural Language Processing (NLP). Efforts such as ACL SIGARAB [^1] https://www.sigarab.org/ have increased visibility and coordination across the community, but the development of robust, domain-specific Arabic language models and datasets continues to lag behind progress in high-resource languages such as English and Chinese [el2010using].
This gap is driven by persistent data scarcity and the linguistic characteristics of Arabic itself. The language is diglossic, spanning Classical Arabic, Modern Standard Arabic (MSA), and a wide range of regional dialects that vary lexically, morphologically, and syntactically. In addition, its right-to-left script, lack of capitalisation, and frequent omission of diacritics—short vowel markers that help disambiguate homographs—create further difficulties for tokenisation, parsing, and semantic interpretation [el-haj-rayson-2016-osman].
Recent advances in Arabic NLP—through models such as AraBERT [antoun_arabert_2021], ARBERT, MARBERT [abdul-mageed_arbert_2021], and generative models like AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022]—have substantially improved general text understanding. However, these models are typically trained on broad web or news data and therefore lack consistent exposure to specialised financial or technical language. By contrast, English NLP has benefited from domain-adapted architectures such as FinBERT [yang2020finbert] and BioBERT [lee2020biobert], which have demonstrated clear gains in financial and biomedical tasks. The absence of similar domain adaptation efforts for Arabic underscores the need for targeted pretraining strategies capable of capturing the linguistic and conceptual characteristics of financial discourse.
This need is amplified by the rapid expansion of the Arab financial sector over the past two decades, with markets such as Tadawul [^2] https://www.saudiexchange.sa, the Dubai Financial Market (DFM) [^3] https://www.dfm.ae, and the Abu Dhabi Securities Exchange (ADX) [^4] https://www.adx.ae now ranking among the world’s most active [zmandar2021joint]. The resulting increase in Arabic financial narratives—corporate disclosures, regulatory filings, market analyses, and news reports—has produced texts that are long, information-dense, and stylistically formal, making them well suited to automatic summarisation [zmandar2021multilingual]. Yet most Arabic summarisation systems continue to focus on general news, leaving financial text underexplored despite its demand for accuracy, numerical precision, and stylistic clarity [el2020financial, zmandar2021financial, zavitsanos2023financial].
@@ II Related Work
Text summarisation has long been a core task in Natural Language Processing (NLP), particularly within journalism and information retrieval. Early systems were largely extractive, relying on sentence ranking and statistical heuristics such as term frequency, cue words, and sentence position [filippova2009company, berger2000ocelot]. Although these methods captured surface-level salience, they struggled to produce coherent, concise outputs and often generated redundant summaries.
@@
The shift to neural encoder–decoder architectures marked a turning point. Recurrent models with attention mechanisms enabled systems to generate novel text rather than simply selecting existing sentences. One of the earliest demonstrations of fluent headline generation using LSTM-based architectures was presented in [lopyrev2015generating]. Subsequent improvements incorporated coverage mechanisms and reinforcement learning to reduce redundancy and improve factual consistency [see-etal-2017-get, cheng-etal-2016-neural].
@@
Transformer-based models such as BART and T5 further advanced abstractive summarisation. Their text-to-text formulation supported strong generalisation across languages and tasks [raffel_exploring_2020], while multilingual variants like mT5 [xue-etal-2021-mt5] extended these capabilities to low-resource settings through cross-lingual transfer. However, these broadly pretrained models often struggle in specialised domains—such as finance—where numerical reasoning, formal style, and domain-specific terminology require more targeted adaptation.
@@
Recent developments in Arabic NLP have mirrored these global trends. Large-scale pretrained models such as AraBERT [antoun_arabert_2021], ARBERT and MARBERT [abdul-mageed_arbert_2021] have become standard baselines for classification and sentiment analysis, addressing challenges related to rich morphology and dialectal variation. Generative models including AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022], and more recent efforts such as Jasmine [abdul2023jasmine], have extended these capabilities to summarisation and text generation tasks.
@@
A growing body of research now focuses specifically on Arabic abstractive summarisation. [kahla2023fine] demonstrated the benefits of fine-tuning multilingual transformers for Arabic news summarisation, while [morsi2025arabic] proposed transformer-based architectures that improved coherence and fluency over earlier systems. Other work has integrated linguistic information or auxiliary objectives- for example, [essa2025enhanced] combined named entity recognition with summarisation to improve factual preservation, and [elsaid2025hybrid] developed a hybrid RNN–Transformer model to enhance grammatical consistency. Collectively, these studies highlight the feasibility of generative approaches for Arabic, though most efforts remain centred on general news rather than specialised professional domains.
Despite the progress, current Arabic language models are still predominantly general-purpose. Models such as AraBERT, ARBERT, MARBERT, AraBART, and AraT5 are trained on broad web or news corpora and therefore have limited exposure to financial or technical text. A notable exception is FinAraT5 [zmandar2023finarat5], a domain-adapted variant further pretrained on large Arabic financial corpora. However, such efforts remain rare compared with English NLP, where models like FinBERT [yang2020finbert] and BioBERT [lee2020biobert] demonstrate the effectiveness of targeted pretraining for specialised tasks. The scarcity of comparable domain-specific resources for Arabic reinforces the need for focused pretraining and systematic evaluation in financial summarisation.
@@ II-A Datasets for Arabic Summarisation and Financial Corpora
Research in Arabic summarisation has been supported by a range of corpora covering both extractive and abstractive paradigms. Early resources such as Kalimat [ElHaj2013KALIMATAM], a 20,000-article Omani newspaper corpus, and the Essex Arabic Summaries Corpus (EASC) [el2010using], which includes 153 news articles and 765 human-written summaries, provided the foundation for extractive approaches. More recent multilingual datasets have enabled progress in neural and cross-lingual methods. The Arabic Gigaword corpus [Parker2019gigawords], although large and widely used, lacks aligned document–summary pairs and remains commercially restricted. In contrast, openly available abstractive benchmarks such as XLSum [hasan-etal-2021-xl] and WikiLingua [ladhak-etal-2020-wikilingua] introduced approximately 40,000 and 21,000 Arabic examples respectively, though their content diverges from financial and professional prose.
Additional Arabic-specific datasets have further expanded training resources. SumArabic [SumArabic] provides 84,764 high-quality abstractive text–summary pairs built from Common Crawl data and has been used to benchmark RNN-based and Transformer-based models. HASD [HASD] introduces a 43,000-article benchmark containing both extractive and abstractive summaries, alongside an additional 150,000-sample abstractive dataset (AASD) and a modified version of EASC with new abstractive annotations. While these datasets have accelerated progress in Arabic abstractive summarisation, they target general content and do not capture the stylistic and numerical characteristics of financial narratives.
@@
Earlier work on Arabic summarisation relied heavily on extractive methods such as TF–IDF ranking, lexical scoring, and graph-based centrality measures [El-Haj-2011, Douzidia2004LakhasAA, el2013using, koulali2013arabic]. More recent studies have shifted towards neural models, with multilingual architectures such as mBART and mT5 fine-tuned for Arabic summarisation [kahla-etal-2021-cross], yielding improvements in fluency and abstraction.
@@
However, unlike English, Arabic still lacks large, curated corpora for specialised domains. Resources such as BORSAH [ALSHAHRANI18] and the Arabic Business and Management Corpus (ABMC) [^5] https://elhaj.uk/corpora.html offer valuable financial and economic text collections, but their primary use has been sentiment analysis and classification rather than summarisation. As a result, progress in Arabic abstractive summarisation—particularly for financial text—has been limited by the absence of domain-focused, large-scale datasets.
Despite this progress, existing resources still fail to meet the requirements of Arabic financial summarisation. Most available datasets focus on general news or multilingual content, leaving domain-specific financial material largely unaddressed. The only notable effort in this direction is the FinAraSum dataset [zmandar2023finarat5], which contains approximately 44,900 financial news articles collected from the Al Arabiya Aswaq website over a decade and was designed for headline generation. While valuable, FinAraSum is limited in size, covers a narrower range of financial reporting, and—crucially—has not been released publicly, restricting reproducibility and wider use. As a result, researchers lack an open, large-scale Arabic financial summarisation benchmark, leaving a significant gap in resources for training, evaluation, and domain-adapted modelling.
The AraFinNews dataset introduced in this paper fills this gap by providing over 200,000 high-quality Arabic financial article–headline pairs, enabling research in summarisation and broader financial NLP.
@@
This unmet intersection between Arabic summarisation and financial text processing presents a significant research opportunity: building datasets and models that can handle numerically dense, terminologically specialised, and stylistically formal financial narratives. Addressing this need is essential for developing practical, context-aware summarisation systems for professional Arabic applications.
@@ III AraFinNews: The Arabic Financial News Dataset
To address the lack of large, domain-specific resources for Arabic summarisation, we introduce AraFinNews, the largest publicly available Arabic financial news dataset to date. The corpus comprises 212,500 article–headline pairs sourced through Argaam.com, a major financial media platform in the Gulf region that aggregates material from several well-established outlets, including world financial news from the Financial Times, Forbes, the World Gold Council and Bloomberg. Each headline serves as an abstractive summary of its corresponding article, reflecting the concise, information-rich style characteristic of professional Arabic financial journalism. The dataset spans corporate disclosures, market analyses and macroeconomic reports, offering a solid benchmark for abstractive summarisation as well as related NLP tasks such as named entity recognition, event extraction and sentiment analysis.
AraFinNews covers a continuous period of nearly ten years, from 8 October 2015 to 29 July 2025, capturing key economic cycles, market events, and policy developments across the world. This long-term temporal span ensures the dataset reflects evolving financial language, trends, and reporting styles over time—making it a valuable resource for both synchronic and diachronic analyses of Arabic financial discourse.
@@ IV Financial Large Language Models for Arabic Summarisation
To address the limitations of general-purpose Arabic models, recent work introduced Arabic-adapted transformers such as AraBERT [antoun2020arabert] and AraT5 [nagoudi_arat5_2022], as well as the domain-specialised FinAraT5 [zmandar2023finarat5] [^7] https://huggingface.co/drelhaj/FinAraT5. In this study, we evaluate these models—alongside multilingual baselines—under a unified experimental setup to assess their ability to generate coherent financial headlines for AraFinNews. The following subsections detail the model selection, fine-tuning configuration, and evaluation framework adopted in this work.
@@ IV-A Model Selection and Comparative Setup
All models adopt a text-to-text paradigm supporting multiple generation tasks through consistent input–output formatting. Each model was fine-tuned on the AraFinNews training split using identical hyperparameters to ensure comparability, with the development set used for validation and the held-out test set reserved exclusively for final evaluation. Fine-tuning employed the Hugging Face Transformers library with the AdamW optimiser, a learning rate of $5\times 10^{-5}$ with linear warm-up, an effective batch size of 32 achieved via gradient accumulation over four steps, and a maximum sequence length of 512 tokens. Evaluation assesses how domain adaptation influences summarisation performance.
@@ IV-B Arabic Financial Summarisation Pipeline
Our summarisation pipeline integrates FinAraT5 [zmandar2023finarat5] as the core generation component for producing abstractive financial headlines. The model operates within a sequence-to-sequence framework based on AraT5 [nagoudi_arat5_2022], following the text-to-text paradigm of T5 [raffel_exploring_2020]. Within this setup, full-length financial news articles are transformed into concise headline-style summaries that reflect the conventions of Modern Standard Arabic financial reporting.
Figure 3 illustrates the overall workflow, in which FinAraT5 receives the processed article text as input and generates an abstractive headline output. This component forms the central step of the pipeline used to evaluate summarisation quality on AraFinNews.
@@
FinAraT5 was pretrained using the span-corruption denoising objective introduced in the original T5 framework [raffel_exploring_2020], where contiguous spans of text are randomly masked and the model learns to reconstruct the missing content. This self-supervised objective enables large-scale learning without labelled data and captures long-range contextual dependencies.
Building on the methodology of AraT5 [nagoudi_arat5_2022], FinAraT5 extends this pretraining through continued exposure to raw Arabic financial documents, including large-scale news sources such as AraFinNews. This additional phase enriches the model’s handling of financial terminology, numerical expressions, and the discourse structures characteristic of economic reporting. Figure 4 illustrates the span-corruption task used during pretraining.
@@
FinAraT5 was trained using the original TensorFlow implementation of T5888https://github.com/google-research/text-to-text-transfer-transformer and managed via the t5 library [^9] https://pypi.org/project/t5/. As the model predates later frameworks such as T5X, distributed training was performed using Mesh TensorFlow.
This pretraining strategy provides a strong foundation for downstream generative tasks, particularly abstractive summarisation. Continued exposure to financial text improves the model’s treatment of specialised vocabulary and high-frequency expressions such as (“stocks, profits, indices, returns”), offering advantages over general-purpose Arabic models like AraT5 and mT5. The fine-tuning configuration used in this study is summarised in Table V.